MIDC News
Year of 2007
December 12, 2007
- 11:34 MST: Made an adjustment to SRRL BMS valid quality range for snow depth at 11:28 MST today. -Afshin Andreas
- 09:35 MST: The SRRL BMS logger program was updated yesterday @ 14:00 MST, to improve the data processing of the snow depth based on the depth quality indicator. This will help in automatically filtering bogus snow depth values without having to look at the quality indicator. -Afshin Andreas
- 13:16 MST: Data from the Oak Ridge National Laboratory (ORNL) Rotating Shadowband Radiometer (RSR2) is now available on MIDC. Data is currently updated every 2 hours and data collection started on September 12, 2007. The station also includes a meteorological setup.
-Afshin Andreas
- 14:54 MST: Problem with SRRL LICOR Global Horiz Spectrometer is due to faulty RS232 communications cable, which will hopefully be fixed in the near future. -Afshin Andreas
- 14:55 MST: The SRRL BMS CR23X logger that collects data from instruments on the deck, was swapped out with another logger from 14:19 to 14:39 MST today. Various instruments could read missing or bogus during that time. -Afshin Andreas
- 07:17 MST: There is intermittent communcations problems with the SRRL Global LI-1800 spectrometer. -Afshin Andreas
- 08:16 MST: SRRL Global Horiz Spectral data collection has resumed today as of 08:10 MST. Problem was caused by unknown RS232 communications problem. -Afshin Andreas
- 14:03 MST: Communications with the Global LI-1800 was lost at 12:06 MST, problem unknown. -Afshin Andreas
- 11:32 MST: The SRRL Global Horizontal LI-1800 Spectroradiometer was removed from service @ 9:22 MST, calibrated against NIST F597, and returned to service @ 11:28 MST today. -Afshin Andreas
- 10:46 MST: A "Preset calculation" of "Wind Direction Delta" is now available for SRRL BMS and NWTC M2 live/daily data/plots. This can be used instead of the "User-defined calculation" to properly compute the difference between the wind direction at different levels. -Afshin Andreas
- 09:28 MST: The SRRL BMS Diffuse Shadowband PSP is working again, problem unknown. -Afshin Andreas
- 07:41 MST: The SRRL Rotatating Shadowband Pyranometers (both ATI and RSP2), were upgraded from a modem to an NL100, yesterday. Both sites now update every 30 minutes (instead of every hour). -Afshin Andreas
- 07:27 MST: There is a problem with the SRRL BMS Diffuse Shadowband PSP, which stared today @ 07:22 MST. Problem will be investigated. -Afshin Andreas
- 08:31 MST: August 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 08:24 MST: There is a problem with the SRRL BMS Global PYR-P that started on August 14. Problem will be investigated. -Afshin Andreas
- 08:12 MST: August 2007 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 12:51 MST: The SRRL BMS Research MT0 and MT1 field outputs, were replaced at 11:51 MST with a Snow Depth Quality indicator and a met. tower CR10X table overun counter. -Afshin Andreas
- 12:42 MST: The SRRL BMS Senix "Snow Depth" sensor was removed from service at 10:13 MST today. It was replaced with a Campbell Scientific SR50A S/N 1263 Snow Depth sensor at 11:57 MST today. This sensor is designed for snow and should provide more accurate readings. -Afshin Andreas
- 12:33 MST: The SRRL BMS Tower HMP45C, which measured Dry Bulb Temperature and Relative Humidity, was replaced with a new sensor (s/n C1920073, factory cal 5-9-07) today at 11:11 MST. -Afshin Andreas
- 07:28 MST: The Nevada Power Clark Station NIP was replaced with calibrated instrument on Aug 9 @ approx 9:00 PST. The Global Pyranometer was replaced on Aug 9 @ approx 11:00 PST. The calibration factors were updated in the logger program on Aug 10 @ 6:25 PST. -Afshin Andreas
- 13:37 MST: July 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 13:04 MST: July 2007 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 12:54 MST: The NWTC M2 Global PSP was swapped out with a calibrated instrument today @ 9:52 MST. Calibration factor was updated in logger today @ 11:46 MST. -Afshin Andreas
- 11:23 MST: The Bird Estimated Direct for the SRRL and XECS RSP v2 has been corrected. -Afshin Andreas
- 10:06 MST: June 2007 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 09:43 MST: June 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 12:33 MST: The SRRL BMS EKO MS-110 photometer which had been removed from on 06/16/2005 was officially removed from the current BMS datastream configuration today. The 8 freed channels have been re-asigned to "Research Channels". -Afshin Andreas
- 14:16 MST: The Bird Estimated Direct Normal was added to the SRRL Irradiance Inc RSP2 at 14:05 MST today. Note: There may be a problem with this calculation. -Afshin Andreas
- 09:03 MST: As of July 6, 2007, live graphical displays are available for the following sites: Xcel Energy Comanche Station, Nevada Power Clark Station, University of Nevada, Las Vegas, Elizabeth City State University. -Afshin Andreas
- 15:12 MST: An Error Flag was added to the SRRL Irradiance RSP v2 data stream today @ 14:57 MST. -Afshin Andreas
- 08:07 MST: The shading arm on the SRRL BMS Sci-Tec tracker was bent out of position starting on 7/7 at sunrise to 7/9/2007 at 7:00 MST. Affected instruments were: Diffuse 8-48 and PSP and the Direct CH1. Also, the Diffuse UVB-1 was running without temperature control during that same time, due to another tracker tripping a GFI. -Afshin Andreas
- 12:05 MST: The data collection frequency for Elizabeth City State University was increased from 3 hours to 1 hour. -Afshin Andreas
- 12:04 MST: The Xcel Energy Comanche Station datalogger program was updated @ 8:07 MST today. The following outputs were added to the data stream: uncorrected global/direct/diffuse, peak wind speed, error flag, and accumulated precipitation. -Afshin Andreas
- 14:46 MST: The following SRRL BMS Global instruments (both PSP & LI-200) were swapped out today with instruments that were calibrated in BORCAL 2007-02: 90-North, 90-West, and Upwelling. Logger was updated with new cal factors today @ 14:06 MST. -Afshin Andreas
- 11:21 MST: The Xcel Energy Comanche Station was added to the MIDC on June 29. Data collection started on June 20 and is currently being updated every 15 minutes. The data outputs from this site is currently under construction (a few fields will be added). -Afshin Andreas
- 13:24 MST: The SRRL BSRN Direct NIP and Diffuse PSP were replaced with instruments calibrated in BORCAL 2007-01 & 02 today. New cal factors were entered into the logger @ 13:01 MST today. -Afshin Andreas
- 10:15 MST: The following SRRL BMS instruments were swapped out today with instruments which were calibrated in BORCAL 2007-01 & 02: Global & Direct RG780, Diffuse PSP (Shadowband), Diffuse PSP (ventilated), Diffuse 8-48. New cal factors for these instruments as well as the Global CM22 were entered into the logger @ 9:11 MST today. -Afshin Andreas
- 10:12 MST: The SRRL BMS Global Normal CM21 was swapped out with an instrument which was calibrated in BORCAL 2007-01 today. The Global CM3 (on CNR1) was calibrated using BORCAL 2007-01 reference data. New cal factors were entered in the logger @ 8:01 MST today. -Afshin Andreas
- 09:18 MST: On June 19, 2007, the SRRL Schott RSP was shutdown and upgraded by Irradiance Inc. The SRRL Irradiance Inc. RSP v2 data is now available, data collection started June 19 @ 13:13 MST. The instrument will require calibration. -Afshin Andreas
- 08:12 MST: A Handar 435A sensor was installed at South Park Mountain Data (SPMD). As of 07:40 MST today, relative humidity and a secondary temperature is now available. -Afshin Andreas
- 14:33 MST: May 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 14:22 MST: May 2007 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 09:47 MST: The following SRRL BMS instruments were calibrated in BORCAL 2007-01: Global Horizontals: CM6b, PSP (ventilated), PSP, TSP-700, TSP-1, SP-Lite, PYR-P, LI-200; Direct Normals: CH1, NIP #2, NIP #1, LI-201; Diffuse CM22. Logger program was updated at 8:57 MST today with new cal factors. -Afshin Andreas
- 09:44 MST: The SRRL BMS Global CM22 and Diffuse CM22 instruments were swapped between 8:45 to 9:15 MST today. The cal factors were updated in the logger at 8:57 MST. -Afshin Andreas
- 07:40 MST: The SRRL BSRN Global PSP was replaced with an instrument that was calibrated in BORCAL 2007-01. Logger was updated with new cal factor on 6-12-07 @ 13:04 MST. -Afshin Andreas
- 07:39 MST: The following SRRL BMS Global Tilted instruments (both PSP & LI-200) were swapped out with instruments that were calibrated in BORCAL 2007-01: 40-South, 90-East, and 90-South. Logger was updated with new cal factors on 6-12-07 @ 12:33 MST. -Afshin Andreas
- 10:48 MST: The NWTC M2 tower was dropped again today, so that the Wind Speed sensors could be replaced with instruments that have a more recent calibration. Logger program was updated at 10:44 MST today. -Afshin Andreas
- 07:39 MST: Fixed a bug with generation of hourly data (using live or daily high resolution data) that occured with older style data logger output (24:00 timestamp output rather than 0:00), which caused a second hour 0 output. -Afshin Andreas
- 12:20 MST: April 2007 Monthly Data Sets & Reports for ECSU has been generated. -Afshin Andreas
- 12:15 MST: April 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 12:38 MST: The historical global, direct, and diffuse data from NPCS (mentioned in May 16, 2007 news) has been corrected. -Afshin Andreas
- 11:15 MST: The NWTC M2 tower was dropped yesterday for instrument change out. All Wind Speed, Direction, Temperature, and Dew Point sensors were replaced. New calibration constants for Wind Speed and Temperature were entered into the datalogger today @ 9:39 MST. The M2 instrument page has been updated to reflect the change out. Note: The Wind Direction at 2m and 20m is currently not functioning. -Afshin Andreas
- 13:25 MST: For calibration purposes at Nevada Power Clark Station (NPCS) the following instruments were swapped out: Direct on 4/26/2007 11:07 PST and Global on 5/10/2007 at 15:06 PST. New calibration factors were entered in the logger on May 16 @ 12:11 PST. Historical data will be corrected tomorrow. -Afshin Andreas
- 07:25 MST: There is a problem with the SRRL BMS Global PSP (ventilated) which started on May 13 @ ~14:30 MST. Instrument is reading way to high. -Afshin Andreas
- 07:23 MST: Data collection at Bluefield State College resumed May 2, 2007 @ 13:15 EST. The instruments had been disconnected since November 18, 2006 during roof replacement. -Afshin Andreas
- 12:25 MST: March 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 12:12 MST: March 2007 Monthly Data Sets & Reports for ECSU has been generated. -Afshin Andreas
- 13:46 MST: A new shadow band encoder for the SRRL Total Sky Imager TSI-880 was installed today @ 13:37 MST. This will hopefully correct the weekly alignment problem which required a manual reset. -Afshin Andreas
- 11:35 MST: The SRRL Global Horizontal LI-1800 Spectroradiometer was removed from service @ 8:52 MST, calibrated against NIST F597, and returned to service @ 10:38 MST today. -Afshin Andreas
- 11:00 MST: Historical data for Phoenix Federal Correctional Institution (PFCI) from September 25, 2002 to October 31, 2004 has been added to MIDC. Data during that time was 10-minute averages. -Afshin Andreas
- 08:27 MST: As of March 20, the SRRL BMS Instantaneous data collection has been increased from 6-second to 3-second intervals. This affects the following: SRRL BMS Real-Time Weather Display (can only be effectively seen on SRRL building display monitor), Instantaneous Display (Text Based), and Instantaneous ASCII Data Download. These products are only available internal. -Afshin Andreas
- 08:03 MST: February 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 07:30 MST: February 2007 Monthly Data Sets & Reports for ECSU has been generated. -Afshin Andreas
- 13:05 MST: The ECSU data from Feb 21, 22, & 23, 2007 was corrected today for the clock drift that occured on those 3 days. -Afshin Andreas
- 13:20 MST: The 2 minute timeout for external users when requesting a large custom ZIP file (from daily, monthly, site comparison, or solar position) has been fixed using an undisclosed work around. -Afshin Andreas
- 14:28 MST: The NREL Proxy server was recently reconfigured with a two minute timeout. MIDC users could be affected by this if they request a large custom ZIP file (from daily, monthly, site comparison, or solar position) that takes longer than two minutes to create. The result will be an Error 502 Bad Gateway after 2 minutes have elapsed. Internal NREL users can avoid this problem by using the internal MIDC server. External users will have to select a smaller time range, until a solution is found. -Afshin Andreas
- 10:32 MST: The ECSU datalogger clock was fell behind 4 hours and 20 minutes on 2/21/07. Clock was reset today at 12:25 EST. Data will be corrected sometime within 30 days. -Afshin Andreas
- 15:03 MST: January 2007 Monthly Data Sets & Reports for ECSU has been generated. -Afshin Andreas
- 14:59 MST: January 2007 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 11:16 MST: The NWTC M2 data logger program was updated today @ 11:07 MST, a limit was put on Dew Point such that it does not exceed Temp @ 2m. -Afshin Andreas
- 09:46 MST: December 2006 Monthly Data Sets & Reports for ECSU has been generated. -Afshin Andreas
- 09:33 MST: December 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 10:57 MST: Due to communication issues created with a recent upgrade of LoggerNet and the aging SRRL BMS CR23X datalogger. The CR23X had to swapped out with a spare unit on January 12 from 13:18 to 13:35 MST. Several instrument will read missing or bogus during that time. -Afshin Andreas
- 08:12 MST: The was another power outage the NWTC M2 shed last night from 17:23 to 19:58 MST. A new UPS or battery still needs to be purchased. -Afshin Andreas
- 16:06 MST: As of 14:58 MST today, the MIDC server is now duplicated to a new public server. NREL users can still access the internal MIDC server, which is required for special data sets labeled as "NREL or Internal Only". Note: There could be up to a 1 minute delay before data appears on the public server, therefore internal NREL users should use the internal server if quicker data access is desired. -Afshin Andreas
- 14:22 MST: Communcation to the NWTC M2 datalogger was re-estabilished yesterday @ 17:47 MST. There was a power outage and the building's UPS had failed. -Afshin Andreas
- 11:50 MST: The wireless ethernet to the NWTC M2 shed is currently down, last data collection was on January 8, 2007 at 03:29 MST.
-Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]